OpenStudy (anonymous):

y+5 -6y=17-2y? will give medal and fan

OpenStudy (anonymous):

you want the answer or help to get it?

OpenStudy (anonymous):

y+5-6y=17-2y y+5 -(6y-6y)=17-(2y+6y) y+5 =17-8y y+ (5-5)+17+5 y=22

OpenStudy (anonymous):

i got something different @katenafighter

OpenStudy (anonymous):

orly winner gets medal..........GO

OpenStudy (anonymous):

you have to add 6y to each side no subtract @katenafighter

OpenStudy (anonymous):

y=-4

OpenStudy (anonymous):

y+(5-5)=17+5 y+22

OpenStudy (anonymous):

y=5-6y=17-2y -5y+5=17-2y -3y+5=17 -3y=12 y=-4

OpenStudy (anonymous):

*y+5-6y=17-27

OpenStudy (anonymous):

I think your right @shayright1090

OpenStudy (anonymous):

thank you